After cleaning EI source and replacing the two filaments, the autotune report (joined to this post) says that there is a problem on Q1. Any ideas about this problem's origin? Thanks a lot in advance,
An Agilent technician came for the annual maintenance and solved the problem. The collision cell was damaged, he changed it and the autotune went back to normal.
Thanks for your response
i think you've got a small leak in your system: Air/Water Check: H20 0.63% (<=20.00%), O2 5.14% (<=2.50%), N2 19.15% (<=10.00%)
So check your capillary nut at the transferline and try tuning again.
If it worked before you cleaned the source and replaced the filaments, it is possible that you introduced a problem during that procedure. I would start by venting, removing the source, and doing an inspection. Take the source back apart and reassemble it. Are all the lenses in place correctly? Are the insulators in place? Are all the wires connected to the right places? The brown extractor lens wire? Is the source seated all the way? Is the transferline tip in place? How far does the column extend from the transferline?
The Air/Water check is invalid as the spectrum is not good.
And let us know what you find.
It was the collision cell that was damaged, we don't how it happened but the autotune work now
Thanks for your help,
I'm going to try to see that collision cell. Thanks for the update.
Retrieving data ...